# Using LLMs in Generating Design Rationale for Software Architecture Decisions

### v1

@@ 4.3.1. For Accuracy Metrics
In Table 7, LLM-based agents generate more arguments than zero-shot but fewer than the CoT method, as Aspect_Identifier agent is required to specify up to six most relevant perspectives for architecture analysis. The DR generated by LLM-based agents has a lower average Precision than the zero-shot method but higher than the CoT method, while its F1-score is comparable to the zero-shot method and superior to the CoT method. LLM-based agents show a notable improvement in Recall, with an increase of 19.8% compared to the zero-shot method and 4.5% compared to the CoT method. This result indicates that LLM-based agents can identify and focus on the key factors that align with human experts’ arguments in architecture decisions.
@@ 4.3.2. For IHUM-Category
The proportion of “Uncertain” and “Misleading” arguments in the DR generated by LLM-based agents has decreased compared to other prompting strategies. The decline in “Uncertain” arguments is relatively marginal, while the reduction in “Misleading” arguments is more significant, with the proportion decreasing to 1.59%. According to the output of each agent, the reduction in “Misleading” arguments is attributed to Aspect_Reviewer and Trade-off_Analyst agents. Aspect_Reviewer agent plays a critical role in identifying and rectifying errors within the arguments. For example, in the DR generated by gpt-3.5-turbo for Discussion #787, Aspect_Reviewer agent highlights that the performance analysis by Aspect_Analyst lacks sufficient rigor, emphasizing that factors such as concurrency, resource management, and scalability are also crucial in evaluating whether the architecture decision would lead to performance improvements. This analysis of Aspect_Reviewer agent effectively prevents the inclusion of vague and inadequately substantiated conclusions of “enhancing system efficiency” as a reference in the final trade-off analysis. Therefore, in the final DR, the performance discussion focuses on potential concurrency issues in a multi-threaded environment and the need for effective resource management, rather than claiming a definitive performance improvement. Trade-off_Analyst agent is required to conduct a trade-off analysis according to the analysis from different aspects and provide the final DR. We observed that the DR generated by the multi-agent system based on five LLMs incorporates a consideration of both positive and negative impacts, which is similar to the DR generated with the CoT template.
Key findings of RQ3 Compared to the zero-shot and CoT methods, the DR generated by LLM-based agents includes the most arguments align with human experts, exhibiting the highest Recall. The Precision is slightly lower than that of the zero-shot method due to the increased arguments number, while the F1-score is comparable to the zero-shot method. The increase in “Insightful” arguments and the decrease in “Helpful” arguments ensure that the DR generated by LLM-based agents provides a comprehensive perspective on the architectural decision while minimizing the inclusion of low-relevance arguments. The DR generated by LLM-based agents includes the fewest “Uncertain” and “Misleading” arguments, primarily due to the modifications made by Aspect_Reviewer agent and the analysis provided by Trade-off_Analyst agent.
@@ 5.1. Pros & Cons of Three Prompting Strategies
Zero-shot can generate effective DR with single-round dialogue, however, its key limitation lies in the lack of comprehensive analysis and the inclusion of some irrelevant content.
@@
Zero-shot is a basic prompting strategy that provides LLMs with the context of architecture decisions and basic instructions to generate DR, requiring only one single LLM invocation. According to the average performance presented in Table 7, the DR generated with zero-shot method demonstrates the highest Precision (0.278) and an F1-score (0.381) comparable to that of LLM-based agents. For other metrics, the zero-shot approach also shows no significant gap compared to other methods. The performance of the zero-shot method indicates that LLMs are capable of generating DR for architecture decisions without additional training.
@@
However, the DR generated through zero-shot rarely identifies the potential issues of architecture decisions for analysis. For the five LLMs, only 16% of the DR generated by the zero-shot method includes an analysis of architectural weaknesses. Therefore, the arguments provided in the DR generated by zero-shot may lack comprehensive analysis. For instance, the DR generated for SO #77745340 by gpt-3.5-turbo incorrectly judges the performance of the architecture decision due to overlooked IPC overhead, as mentioned in Section 4.2.2. The lack of analysis of potential architectural weaknesses may result in skewed conclusions in arguments, which contributes to the highest proportion of “Misleading” arguments in the DR generated with the zero-shot method. In addition, the prompt template for the zero-shot method does not specify any requirements for the output. As a result, the analysis of DR generated by the zero-shot method is unpredictable and may occasionally include content unrelated to the DR. For example, the DR generated by gemini-1.0-pro for Issue #45 also suggests other architecture decisions to mitigate “thundering-herd scenarios”, where multiple clients or processes simultaneously compete for a shared resource. Such information in the output is unrelated to the DR and may not align with the expectations of software architects regarding the generated content.
[…]
@@
CoT enhances the comprehensiveness of DR by incorporating trade-off analysis compared to the zero-shot method, however, it lacks in-depth analysis for individual arguments.
The CoT method instructs LLMs to generate DR step by step according to the provided template. All five LLMs correctly follow the instructions in the CoT template during DR generation, outlining the potential advantages and disadvantages of architecture decisions for the trade-off analysis in the final step. Compared to the zero-shot method, the CoT method enhances the reasoning capabilities in DR generation, enabling more thorough analysis during the trade-off process and reducing the proportion of “Misleading” arguments compared to the zero-shot method, as explained in Section 4.2.2.
@@
However, some argument points in the DR generated using the CoT templates lacks the analytical depth. According to Table 7, the number of argument points generated by CoT is 26.7% higher than that generated by the zero-shot method and 8.5% higher than that generated by LLM-based agents. Although these arguments provide diverse perspectives on understanding architecture decisions, some of them are merely superficial statements without in-depth analysis and reasoning. For example, in the generated DR for Issue #2378 by gpt-3.5-turbo with the CoT method, the discussions on “increased code complexity” just briefly mention the difficulty of implementing the architecture without a detailed analysis of the specific challenges or potential implications of this added complexity. Similarly, the generated DR also does not provide specific analysis when considering “future reusability”. Therefore, software architects could struggle to derive valuable insights from such argument points due to the lack of in-depth analysis.
[…]
@@
LLM-based agents use multi-agent collaboration for information gathering, reasoning, and verification, however, excessive interactions between agents increase conversation rounds with LLMs, and certain agent functionalities remain to be further optimized.
@@
According to Fig. 5, the reasoning process of LLM-based agents involves two steps: Aspect_Analyst agent analyzes each aspect related to architecture decisions, while Trade-off_Analyst agent provides the final DR based on the analysis of these aspects. By analyzing each architecture-related aspect individually, Aspect_Analyst can provide a more detailed analysis for every aspect, mitigating the issue in CoT where arguments are outlined without sufficient analysis. Trade-off_Analyst agent is tasked with performing a trade-off analysis based on the evaluation of different aspects of architecture decisions, which have been reviewed by Aspect_Reviewer agent beforehand. Therefore, the final DR generated by Trade-off_Analyst agent also includes an analysis of the advantages and disadvantages of architecture decisions, similar to the CoT method, ensuring the comprehensiveness of the generated DR. By leveraging these two layers of reasoning (i.e., Aspect_Reviewer and Trade-off_Analyst agents), LLM-based agents generate DR that better captures key concerns of human experts in architecture decisions than CoT. According to Table 7, the DR generated by LLM-based agents exhibits the highest Recall (0.715) among the three prompting strategies, which indicates that it includes most arguments that align with those of human experts. Besides, the review process executed by Aspect_Reviewer, which checks and modifies the analysis results of each aspect by Aspect_Analyst agent, leads to a decreased proportion of “Misleading” arguments in the generated DR as explained in Section 4.3.2.
@@
However, one of the issues of LLM-based agents in generating DR is the increased conversation rounds with LLMs. The generation of one DR may require up to 20 rounds of conversation with the LLMs, which is significantly greater than required by the zero-shot and CoT methods. In addition, some functions of the agents can be further optimized. For instance, it is challenging to assess the value of external knowledge provided by Information_Collector agent. Although certain DR, such as the one for Issue #1517 generated by mistral-7B, explicitly integrates “recent” findings to support the arguments, the impact of external knowledge is difficult to discern in most cases. Therefore, Aspect_Analyst could incorporate functionality for tracking external knowledge in order to clearly identify which arguments are derived from what external knowledge. Furthermore, compared to human experts, the arguments in the LLM-generated DR typically lack concrete code examples. Therefore, equipping Information_Collector with the ability to automatically retrieve open-source project code or code from Q&A sites relevant to the architecture problem, and enabling Aspect_Analyst to generate context-based code examples, would help LLM-based agents to provide DR closer to practical development scenarios.
@@ 5.2. Strengths and Limitations of LLM-Generated DR
[…]
Strength 1: LLMs are capable of identifying the key factors of DR considered by human experts. During the evaluation process, the arguments of DR provided by human experts are regarded as ground truth to quantitatively evaluate the accuracy metrics of LLM-generated DR. Table 7 shows that LLMs exhibit an average Recall higher than 0.6 with all three prompting strategies, in which LLM-based agents even exceed 0.7. This result reveals that most of the key points considered by human experts when making architecture decisions are identified in LLM-generated DR. To enable LLMs to perform multidimensional analysis from different aspects for each architecture problem, we did not explicitly specify in the prompt which perspectives should be focused on. Therefore, the Recall results suggest that LLMs have the potential to reason the key factors of DR based on the specific context of architecture problems and their corresponding architecture decisions.
[…]
@@
The number of arguments in the LLM-generated DR is significantly higher than that provided by human experts. For each architecture decision, human experts provide an average of 1.98 arguments in DR, while LLMs generate an average of 4.55, 5.77, and 5.31 arguments using the three prompting strategies (i.e., zero-shot, CoT, LLM-based agents), respectively. Despite covering most of the arguments provided by human experts, the greater number of arguments in LLM-generated DR leads to a Precision below 0.3 and an F1-score below 0.4. The result suggests that LLMs typically analyze a broader range of aspects in DR for architecture decisions compared to human experts. Human experts may prioritize the feasibility of architecture implementation and its alignment with the project requirements, potentially overlooking other critical considerations. In SO #77579165, the expert suggested “creating a concrete repository class and mocking it for testing in Riverpod and Flutter”, simplifying implementation by synchronizing two files for efficiency and clear responsibility separation. In contrast, the DR generated by llama3-8B expands on this by further emphasizing the separation of concerns, testability, and reusability. Thus, an additional value of LLM-generated DR is the potential ability to complement the DR provided by human experts and help architects to understand architecture decisions from diverse perspectives.
[…]
@@
While LLMs can generate DR from perspectives aligned with those of human experts, they often lack the depth of practical experience and contextual understanding, which are essential characteristics of human expertise. For example, in Issue #65, “Debian” was chosen as the operating system for the container image as an architecture decision. This decision was explained by human experts who not only emphasized Debian’s advantages, but also considered the time loss associated with code refactoring incurred due to migrations from other Linux distributions (e.g., Amazon Linux 2, Oracle Linux, and Universal Base Image) to Debian-based systems. This practice-based analysis is more persuasive for supporting the architecture decision, whereas the DR generated by LLMs typically lacks such experience-driven arguments. In Discussions #442, while the DR generated by LLMs also addressed the complexity of implementing the architecture at the code level, human experts were able to explicitly identify specific potential technical barriers according to the architecture context. Therefore, enabling LLMs to exhibit the experience-driven insights and contextual sensitivity of human experts when analyzing arguments is crucial to improving the quality of the generated DR.
[…]

### v3

@@ 4.3.1. For Accuracy Metrics
In Table 8, LLM-based agents generate more arguments than zero-shot but fewer than the CoT method, as Aspect_Identifier agent is required to specify up to six most relevant perspectives for architecture analysis. Compared to the zero-shot method, llama3-8B shows the largest increase in the number of generated DR (+1.4), followed by mistral-7B (+0.9) and gemini-1.0-pro (+0.8), while gpt-3.5-turbo (+0.5) and gpt-4-0613 (+0.3) exhibit the smallest increases. The DR generated by LLM-based agents has a lower average Precision than the zero-shot method but higher than the CoT method, while its F1-score is comparable to the zero-shot method and superior to the CoT method. However, the slightly lower Precision of LLM-based agents compared to that of the zero-shot method does not necessarily indicate that the performance of LLM-based agents in generating DR is inferior. In fact, LLM-based agents generated most arguments that align with those provided by human experts, and show a notable improvement in Recall, with an increase of 19.8% compared to the zero-shot method and 4.5% compared to the CoT method. This result indicates that LLM-based agents can identify and focus on the key factors that align with human experts’ arguments in architecture decisions. However, in comparison to the zero-shot method, LLM-based agents generated a larger number of arguments, including many that were not mentioned by human experts, which contributed to a modest 0.01 decrease in Precision. These additional arguments not mentioned by human experts did not necessarily diminish the overall quality of the DR generated by LLM-based agents. In fact, most of the arguments that contributed to the slight drop in Precision were considered “Helpful” for understanding the architecture decision, as detailed in Section 4.3.2.
@@ 4.3.2. For IHUM-Category
The proportion of “Uncertain” and “Misleading” arguments in the DR generated by LLM-based agents has decreased compared to other prompting strategies. The decline in “Uncertain” arguments is relatively marginal, while the reduction in “Misleading” arguments is more significant, with the proportion decreasing to 1.59%. According to the output of each agent, the reduction in “Misleading” arguments is attributed to Aspect_Reviewer and Trade-off_Analyst agents. Aspect_Reviewer agent plays a critical role in identifying and rectifying errors within the arguments. For example, in the DR generated by gpt-3.5-turbo for Discussion #787, Aspect_Reviewer agent highlights that the performance analysis by Aspect_Analyst lacks sufficient rigor, emphasizing that factors such as concurrency, resource management, and scalability are also crucial in evaluating whether the architecture decision would lead to performance improvements. This analysis of Aspect_Reviewer agent effectively prevents the inclusion of vague and inadequately substantiated conclusions of “enhancing system efficiency” as a reference in the final trade-off analysis. Therefore, in the final DR, the performance discussion focuses on potential concurrency issues in a multi-threaded environment and the need for effective resource management, rather than claiming a definitive performance improvement. Trade-off_Analyst agent is required to conduct a trade-off analysis according to the analysis from different aspects and provide the final DR. We observed that the DR generated by the multi-agent system based on five LLMs incorporates a consideration of both positive and negative impacts, which is similar to the DR generated with the CoT template. Furthermore, due to the absence of structural constraints in the Trade-off_Analyst agent, the generated DR exhibits variation in content organization across different models. Notably, gpt-3.5-turbo and gpt-4-0613, tend to perform trade-off analysis within each argument, directly balancing the pros and cons of a specific architectural aspect. In contrast, gemini-1.0-pro, mistral-7B, and llama3-8b tend to list separate arguments, usually outlining advantages and disadvantages, and then perform a summary-level trade-off across them.
Key findings of RQ3 Compared to the zero-shot and CoT methods, the DR generated by LLM-based agents includes the most arguments that align with human experts, exhibiting the highest Recall. The Precision is slightly lower than that of the zero-shot method, while the F1-score is comparable to the zero-shot method. The increase in “Insightful” arguments and the decrease in “Helpful” arguments ensure that the DR generated by LLM-based agents provides a comprehensive perspective on the architecture decision while minimizing the inclusion of low-relevance arguments. The DR generated by LLM-based agents includes the fewest “Uncertain” and “Misleading” arguments, primarily due to the modifications made by Aspect_Reviewer agent and the analysis provided by Trade-off_Analyst agent.
@@
4.4. RQ4: How do practitioners perceive LLM-generated DR of architecture decisions in practice?
[…]
Table 9. Participants’ experience in software development and architecture design
Participant Software Development Experience Architecture Design Experience
P1 3 years 2 years
P2 7 years 2 years
P3 3 years 3 years
P4 25 years 10 years
P5 3 years 2 years
P6 2 years 2 years
  • IQ1.1: Concerning the DR generated by LLMs, two practitioners (P5, P6) chose to “generally trust it”. P2 reported that he could “trust it with some caution”, while P1, P3 and P4 indicated that they could only “trust it with substantial human review”. Overall, all the participants recognized the value of LLM-generated DR, but they also expressed reservations about fully relying on LLMs to recover DR in practice. There are three primary reasons highlighted by the six practitioners for their worries: (1) LLMs are unable to fully understand the requirements of the project that architecture decisions must fulfill (P1, P2, P5). The three practitioners consistently emphasized the complexity of clarifying the project requirements in practice. As a result, even P5 and P6, who demonstrated the highest level of trust in LLM-generated DR, cautioned that “LLMs may fail to fully understand specific project requirements”. (2) Human review constitutes an indispensable step in practice to ensure that architectural solutions are well-reasoned and appropriate (P1, P4). No matter the DR is provided by human experts or generated by LLMs, P4, based on his practical experience, stated that “validation should involve discussion among multiple human experts”. (3) Some arguments in the DR generated by LLMs are not accurate or reliable (P3, P4). From the perspective of P3, the DR generated by LLMs are of “mixed quality”. Therefore, he required more thorough human validation before accepting the generated DR.
[…]
  • IQ2.1: Among the six practitioners, P3 stated that he would definitely use LLMs to generate DR in practice. P1, P5 and P6 indicated that they probably would use LLM-generated DR in practice, while P2 and P4 were not sure about their intentions.
  • IQ2.2: In our interviews, practitioners mentioned four scenarios that are suitable for applying LLM-generated DR in practice: (1) Comparing different architecture design (P1, P3, P6): P1 and P3 considered that the DR generated by LLMs could efficiently help them better identify the key differences and trade-offs between architecture designs by revealing the underlying motivations and objectives behind each design. (2) Understanding unfamiliar architecture that exceeds the team’s prior experience (P1, P3, P4): P1, P3, and P4 concurred that the DR produced by LLMs enables them to rapidly understand architectures that are unfamiliar to them. For instance, P3 noted that the DR produced by LLMs was particularly valuable in aiding comprehension of specific details while studying a new architecture. P4 mentioned that when the company’s business needs to expand into entirely new domains, the DR generated by LLMs can help them quickly identify which architectures are applicable. (3) Gaining deeper insight into architecture decisions made in the current system (P2, P5, P6): According to P2, once the architectural layers are defined, LLM-generated DR can analyze the design from multiple perspectives based on architectural context. This may lead to more well-founded insights and help developers better understand the details of architecture. P5 and P6 believed that LLM-generated DR is especially valuable in the absence of documentation concerning architecture design and its rationale, as it helps architects gain a more comprehensive understanding of the system architecture. (4) Guiding architectural evolution (P4, P5): Extending projects with a new functionality module may require modifications or extensions to the existing architecture. P4 noted that in such a scenario, LLM-generated DR can assist in understanding how to update the existing architecture. P5 further emphasized that LLM-generated DR can help architects assess the impact of architectural changes and ensure compatibility between the updated components and the existing architecture.
[…]
Key findings of RQ4 Practitioners rated their trust level of LLM-generated DR between “generally trust it” and “trust it with substantial human review”. Providing LLMs with project context, architecture and DR from similar projects, and domain-specific knowledge is considered useful in improving the credibility of generated DR. The majority of participants expressed a willingness to use LLMs for generating DR in practice, while two of them remained uncertain. “Comparing different architecture design”, “Learning unfamiliar architecture”, “Gaining deeper insight into current architecture decision” and “Guiding architectural evolution” are the four main scenarios for applying generated DR in practice, as identified by practitioners. “Difficulty in providing context”, “limited understanding of complex requirements”, “concerns about reliability”, and “output constraints” are the main challenges in applying LLM-generated DR, as reported by practitioners.
@@ 5.1. Pros & Cons of Three Prompting Strategies
Zero-shot can generate effective DR with single-round dialogue and is commonly employed by practitioners; however, its key limitation lies in the lack of comprehensive analysis and the inclusion of some irrelevant content.
@@
Zero-shot is a basic prompting strategy that provides LLMs with the context of architecture decisions and basic instructions to generate DR, requiring only one single LLM invocation. According to the average performance presented in Table 8, the DR generated with zero-shot method demonstrates the highest Precision (0.278) and an F1-score (0.381) comparable to that of LLM-based agents. For other metrics, the zero-shot approach also shows no significant gap compared to other methods. The performance of the zero-shot method indicates that LLMs are capable of generating DR for architecture decisions without additional training. Besides, practitioners (P2, P3, and P4) emphasized that, when provided with sufficient project context, the zero-shot method can capture fine-grained rationale for architectural issues of specific modules or design concerns. Consequently, the zero-shot method is regarded as practical and widely adopted by practitioners for asking DR in specific architectural contexts.
@@
However, the DR generated through the zero-shot method rarely identifies the potential issues of architecture decisions for analysis. For the five LLMs, only 16% of the DR generated by the zero-shot method includes an analysis of architectural weaknesses. Therefore, the arguments provided in the DR generated by the zero-shot method may lack comprehensive analysis. For instance, the DR generated for SO #77745340 by gpt-3.5-turbo incorrectly judges the performance of the architecture decision due to overlooked IPC overhead, as mentioned in Section 4.2.2. The lack of analysis of potential architectural weaknesses may result in skewed conclusions in arguments, which contributes to the highest proportion of “Misleading” arguments in the DR generated with the zero-shot method. In addition, the prompt template for the zero-shot method does not specify any requirements for the output. As a result, the analysis of DR generated by the zero-shot method is unpredictable and may occasionally include content unrelated to the DR. For example, the DR generated by gemini-1.0-pro for Issue #45 also suggests other architecture decisions to mitigate “thundering-herd scenarios”, where multiple clients or processes simultaneously compete for a shared resource. Such information in the output is unrelated to the DR and may not align with the expectations of software architects regarding the generated content. Therefore, to ensure the DR generated through the zero-shot method aligns with the expected structure, it is recommended to explicitly provide an output template in the prompt during practical use.
@@
CoT enhances the comprehensiveness of DR by incorporating trade-off analysis compared to the zero-shot method, and can potentially accommodate different scenarios; however, it lacks in-depth analysis for individual arguments.
The CoT method instructs LLMs to generate DR step by step according to the provided template. All five LLMs correctly follow the instructions in the CoT template during DR generation, outlining the potential advantages and disadvantages of architecture decisions for the trade-off analysis in the final step. Compared to the zero-shot method, the CoT method enhances the reasoning capabilities in DR generation, enabling more thorough analysis during the trade-off process and reducing the proportion of “Misleading” arguments (2.33%) compared to the zero-shot method (3.24%), as explained in Section 4.2.2. According to the results presented in Table 8, the CoT method achieves higher Recall (0.684) compared to zero-shot (0.627), while its Precision (0.237) and F1-score (0.351) are slightly lower than those of zero-shot and LLM-based agents. Notably, the DR generated by the CoT method contains the highest proportion of “Helpful” arguments (69.42%) among the three approaches, providing architects with richer perspectives for understanding the architecture. Moreover, our experiments demonstrate that the reasoning process embedded in the CoT template can effectively guide LLMs in both reasoning about and generating DR. Therefore, practitioners may tailor the CoT reasoning process by modifying prompt templates to better accommodate different practical scenarios. For instance, “comparing different architecture designs”, in which practitioners consider a valuable scenario, is particularly well-suited to the step-by-step reasoning enabled by the CoT method.
@@
However, some argument points in the DR generated using the CoT templates lack the analytical depth. According to Table 8, the number of argument points generated by CoT is 26.7% higher than that generated by the zero-shot method and 8.5% higher than that generated by LLM-based agents. Although these arguments provide diverse perspectives on understanding architecture decisions, some of them are merely superficial statements without in-depth analysis and reasoning. For example, in the generated DR for Issue #2378 by gpt-3.5-turbo with the CoT method, the discussions on “increased code complexity” just briefly mention the difficulty of implementing the architecture without a detailed analysis of the specific challenges or potential implications of this added complexity. Similarly, the generated DR also does not provide specific analysis when considering “future reusability”. Therefore, software architects could struggle to derive valuable insights from such argument points due to the lack of in-depth analysis.
@@
LLM-based agents use multi-agent collaboration for information gathering, reasoning, and verification; however, excessive interactions between agents increase conversation rounds with LLMs, and certain agent functionalities remain to be further optimized.
@@
LLM-based agents represent the most complex approach among the three prompting strategies. The DR generated by this method achieves the highest Recall (0.715) and F1-score (0.389), while containing the fewest “Uncertain” (4.12%) and “Misleading” arguments (1.59%). Although its Precision is slightly lower than that of the zero-shot method, the higher Recall suggests that LLM-based agents capture a greater number of arguments aligned with those of human experts. According to Fig. 5, the reasoning process of LLM-based agents involves two steps: Aspect_Analyst agent analyzes each aspect related to architecture decisions, while Trade-off_Analyst agent provides the final DR based on the analysis of these aspects. By analyzing each architecture-related aspect individually, Aspect_Analyst can provide a more detailed analysis for every aspect, mitigating the issue in CoT where arguments are outlined without sufficient analysis. Trade-off_Analyst agent is tasked with performing a trade-off analysis based on the evaluation of different aspects of architecture decisions, which have been reviewed by Aspect_Reviewer agent beforehand. Therefore, the final DR generated by Trade-off_Analyst agent also includes an analysis of the advantages and disadvantages of architecture decisions, similar to the CoT method, ensuring the comprehensiveness of the generated DR. By leveraging these two layers of reasoning (i.e., Aspect_Reviewer and Trade-off_Analyst agents), LLM-based agents generate DR that better captures key concerns of human experts in architecture decisions than CoT. According to Table 8, the DR generated by LLM-based agents exhibits the highest Recall (0.715) among the three prompting strategies, which indicates that it includes most arguments that align with those of human experts. Besides, the review process executed by Aspect_Reviewer, which checks and modifies the analysis results of each aspect by Aspect_Analyst agent, leads to a decreased proportion of “Misleading” arguments in the generated DR as explained in Section 4.3.2.
@@
However, one of the issues of LLM-based agents in generating DR is the increased conversation rounds with LLMs. The generation of one DR may require up to 20 rounds of conversation with the LLMs, which is greater than required by the zero-shot and CoT methods. In addition, some functions of the agents can be further optimized. For instance, it is challenging to assess the value of external knowledge provided by Information_Collector agent. Although certain DR, such as the one for Issue #1517 generated by mistral-7B, explicitly integrates “recent findings” to support the arguments, the impact of external knowledge is difficult to discern in most cases. Therefore, Aspect_Analyst could incorporate functionality for tracking external knowledge in order to clearly identify which arguments are derived from what external knowledge. Furthermore, compared to human experts, the arguments in the LLM-generated DR typically lack concrete code examples. Therefore, equipping Information_Collector with the ability to automatically retrieve open-source project code or code from Q&A sites relevant to the architecture problem, and enabling Aspect_Analyst to generate context-based code examples, would help LLM-based agents to provide DR closer to practical development scenarios. Moreover, when applying the LLM-based agent system to a specific project, customizing the capability of Information_Collector to extract architecture-related information in depth from that project is of practical significance, as it alleviates the burden on architects by reducing the need to manually gather project context. This challenge was most frequently highlighted by practitioners for applying LLM-generated DR in practical scenarios.
@@ 5.2. Strengths and Limitations of LLM-Generated DR
Strength 1: LLMs are capable of identifying the key factors of DR considered by human experts.
During the evaluation process, the arguments of DR provided by human experts are regarded as ground truth to quantitatively evaluate the accuracy metrics of LLM-generated DR. Table 8 shows that LLMs exhibit an average Recall higher than 0.6 with all three prompting strategies, in which LLM-based agents even exceed 0.7. This result reveals that most of the key points considered by human experts when making architecture decisions are identified in LLM-generated DR. To enable LLMs to perform multidimensional analysis from different aspects for each architecture problem, we did not explicitly specify in the prompt which perspectives should be focused on. Therefore, the Recall results suggest that LLMs have the potential to reason the key factors of DR based on the specific context of architecture problems and their corresponding architecture decisions.
@@
The number of arguments in the LLM-generated DR is higher than that provided by human experts. For each architecture decision, human experts provide an average of 1.98 arguments in DR, while LLMs generate an average of 4.55, 5.77, and 5.31 arguments using the three prompting strategies (i.e., zero-shot, CoT, LLM-based agents), respectively. Despite covering most of the arguments provided by human experts, the greater number of arguments in LLM-generated DR leads to a Precision below 0.3 and an F1-score below 0.4. Nevertheless, this result does not suggest that the quality of the DR generated by LLMs is low. Accuracy metrics are calculated based on the human experts’ rationale as ground truth; however, these expert-provided DR are not necessarily the only correct answers and may be incomplete. In our IHUM-category evaluation, the majority of arguments that were not aligned with those provided by human experts were still considered “Helpful”. The presence of more than 60% “Helpful” arguments in LLM-generated DR suggests that LLMs typically analyze a broader range of aspects of DR for architecture decisions compared to human experts. Human experts may prioritize the feasibility of architecture implementation and its alignment with the project requirements, potentially overlooking other critical considerations when analyzing DR. For example, in SO #77579165, the expert suggested “creating a concrete repository class and mocking it for testing in Riverpod and Flutter”, simplifying implementation by synchronizing two files for efficiency and clear responsibility separation. In contrast, the DR generated by llama3-8B expands on this by further emphasizing the separation of concerns, testability, and reusability. Thus, an additional value of LLM-generated “Helpful” arguments is to complement the DR provided by human experts and help architects to understand architecture decisions from diverse perspectives. Such advantage was also mentioned by some practitioners during our interviews. For instance, P2 noted that “Compared to practitioners, LLMs are typically able to conduct more comparative analyses from multiple perspectives, thereby providing more reasonable DR”. Besides, some “Helpful” arguments generated by LLMs highlight potential disadvantages of architecture decisions, which reveal trade-offs that were overlooked by human experts. For example, in Discussion #2792, the human expert focused primarily on the benefits of modular design, including improved decoupling, enhanced build efficiency, and the prevention of circular dependencies. In contrast, the DR generated by gpt-3.5-turbo using the LLM-based agents strategy offers a balanced analysis of the pros and cons for each key argument, highlighting both the architectural advantages and the potential overhead associated with management and coordination. When explicitly guided to perform trade-off reasoning through CoT and LLM-based agents strategies, LLM-generated DR can typically reveal architectural trade-offs through those “Helpful” arguments that human experts overlooked.
@@
While LLMs can generate DR from perspectives aligned with those of human experts, they often lack the depth of practical experience and contextual understanding, which are essential characteristics of human expertise. For example, in Issue #65, “Debian” was chosen as the operating system for the container image as an architecture decision. This decision was explained by human experts who not only emphasized Debian’s advantages, but also considered the time loss associated with code refactoring incurred due to migrations from other Linux distributions (e.g., Amazon Linux 2, Oracle Linux, and Universal Base Image) to Debian-based systems. This practice-based analysis is more persuasive for supporting the architecture decision, whereas the DR generated by LLMs typically lacks such experience-driven arguments. In Discussions #442, while the DR generated by LLMs also addressed the complexity of implementing the architecture at the code level, human experts were able to explicitly identify specific potential technical barriers according to the architecture context. Therefore, besides the capability of LLMs in generating DR, the quality of DR is also influenced by the contextual information provided. However, in this study, SO posts usually offer only the necessary context needed to describe a specific architecture problem, without giving a full project background. Although GitHub issues and discussions are associated with specific projects and allow additional background information to be retrieved from the corresponding repositories, the key challenge lies in automatically determining which parts of the context are relevant to a given architecture problem, since manual filtering is highly time-consuming and labor-intensive. Additionally, regarding domain-specific knowledge, the large amount of potentially relevant background information makes it difficult to decide what is most useful for LLMs in generating DR. As a result, determining how to effectively use relevant project context and domain-specific knowledge is crucial for enabling LLMs to exhibit context-aware insights like human experts when analyzing arguments for DR, thereby improving the quality of the generated DR.